OpenStudy (lxelle):

the sum of the first three terms of a geometric series is 64 and the sum of the next three terms is 27. Find the exact value of sum to infinity of the series.

OpenStudy (anonymous):

a1 - first element a1 * q - second a1 * q^2 - third etc so sum of first three: a1 + a1 * q + a1 * q^2 = 64 a1(1+q+q^2) = 64 sum of next three: a1 * q^3 + a1*q^4 + a1*q^5 = 27 a1*q^3(1+q+q^2)=27 can you do it now ?